# From Intention To Implementation: Automating Biomedical Research via LLMs

### v2

@@ 5.3.2 The Effect of LLM-based Filter Agent
To assess the precision of the ratings assigned by the LLM-based filter agent, we engage human reviewers to undertake a parallel assessment using the scoring criteria outlined in Table  1. Both the filter agent and human reviewers evaluate the same set of papers and datasets, with their ratings based solely on each paper’s title and abstract or the dataset’s description. We use Kendall’s W to assess agreement for the ordinal ratings of 150 papers, as it is suitable for measuring concordance in ordinal data. For the binary ratings of 50 datasets, we apply Fleiss’ kappa, which evaluates inter-rater agreement for categorical data among multiple raters. These papers and datasets are sampled from the search results of 10 topics listed in Table 10. Higher values in both metrics indicate stronger agreement, enhancing the reliability of the human ratings. A majority voting mechanism establishes the ground truth for human ratings, which serves as the benchmark for calculating the model’s accuracy. For papers, those with scores of 4 or higher are classified as useful, while others are deemed not useful, and the accuracy is calculated based on this binary classification. The results of this study are represented in Table 5.
@@ 5.4 Performance of Report Generation
To test the impact of our hierarchical report generation method, we compare its reports against those produced by ReAct, Plan-and-Execute, and a naive single-step LLM approach. Specifically, we standardized 20 papers into an experimental report format using each method. None of the methods are equipped with any additional tools. We prompt an LLM to evaluate the reports across four dimensions: logical soundness, level of detail, consistency with the original paper, and readability. The model assigns a score ranging from 1 to 5 for each dimension, with the scoring criteria detailed in Table 6. The final score is calculated as the average across these four dimensions. Figure 9 presents a box plot of the scores for reports generated by the four different methods.
Table 6: Scoring criteria for experimental reports.
Dimension Criterion
@@ 5.5 Performance of Experimental Design
To validate the efficacy of the experimental design module, we construct a comparison experiment against the three baselines introduced in Section 5.1. We employ the search module and Literature Processing module of BioResearcher for 15 research objectives listed in Table 11, thereby obtaining reports and analyses that serve as a knowledge base. The three baselines can invoke search tools to retrieve relevant content from this knowledge base. All experiments are repeated three times, and the evaluation results are averaged and presented in Table 7.
@@
6 Limitations and Future Works
Limitations
In this paper, BioResearcher effectively automates the entire biomedical research process, from literature and dataset searches to experimental design and execution, given a research objective and specified conditions. However, the system does not achieve complete success in executing experiments without manual intervention. This limitation is partly due to the need for further enhancement of the code generator agent’s performance. Additionally, certain anomalies, such as the unavailability of resources specified in the experimental protocols, also hinder fully automated execution. It highlights the necessity of anticipating a broader range of exceptional scenarios and developing corresponding solutions in future work. Moreover, during our practice, we observed that excessively long inputs often result in LLMs producing shorter, more generalized outputs. To enable LLMs to generate detailed and comprehensive experimental protocols, we employed the multi-step, section-wise processing approach at various stages, including experimental report generation and experimental design. However, this iterative method increases the overall cost.
Future Works
To further enhance and expand the capabilities of BioResearcher, we propose two future research directions. Firstly, BioResearcher currently supports only dry lab experiments. Future research could explore integrating automated wet lab technologies, such as Cloud Labs (Arnold, 2022), to extend the system’s applicability to include wet lab studies. Additionally, the system’s utility could be broadened to encompass other scientific disciplines, e.g., computer science, physics, and chemistry. Secondly, hallucination remains a significant challenge in applications based on large language models (LLMs), where the model may generate content that is factually incorrect. In scientific research contexts, this problem can be exacerbated due to the model’s lack of specific domain knowledge. Consequently, addressing LLM hallucination continues to be a crucial area for future research.
@@ 7 Conclusion
Acknowledgments
Yeyun Gong and Chen Lin are the corresponding authors. Chen Lin is supported by the Natural Science Foundation of China (No. 62432011).
Appendix A Details of Baselines Implementations
@@
We implement all baselines using the AutoGen framework Wu et al. (2023) and adopt OpenAI’s text-embedding-ada-002 as the embedding model in RAG applications.
Appendix B Details of Research Objectives
@@
Appendix C Cost Analysis
@@
Appendix D Comparison of protocols generated by 4 systems.

### v4

@@ 5.3.2 The Effect of LLM-based Filter Agent
To assess the precision of the ratings assigned by the LLM-based filter agent, we engage human reviewers to undertake a parallel assessment using the scoring criteria outlined in Table  11. Both the filter agent and human reviewers evaluate the same set of papers and datasets, with their ratings based solely on each paper’s title and abstract or the dataset’s description. We use Kendall’s W to assess agreement for the ordinal ratings of 150 papers, as it is suitable for measuring concordance in ordinal data. For the binary ratings of 50 datasets, we apply Fleiss’ kappa, which evaluates inter-rater agreement for categorical data among multiple raters. These papers and datasets are sampled from the search results of 10 topics listed in Table 9. Higher values in both metrics indicate stronger agreement, enhancing the reliability of the human ratings. A majority voting mechanism establishes the ground truth for human ratings, which serves as the benchmark for calculating the model’s accuracy. For papers, those with scores of 4 or higher are classified as useful, while others are deemed not useful, and the accuracy is calculated based on this binary classification. The results of this study are represented in Table 5.
@@ 5.4 Performance of Report Generation
To test the impact of our hierarchical report generation method, we compare its reports against those produced by ReAct, Plan-and-Execute, and a naive single-step LLM approach. Specifically, we standardized 20 papers into an experimental report format using each method. None of the methods are equipped with any additional tools. We prompt an LLM to evaluate the reports across four dimensions: logical soundness, level of detail, consistency with the original paper, and readability. The model assigns a score ranging from 1 to 5 for each dimension, with the scoring criteria detailed in Appendix D (Table D2). The final score is calculated as the average across these four dimensions. Figure 9 presents a box plot of the scores for reports generated by the four different methods.
@@ 5.5 Performance of Experimental Design
To validate the efficacy of the experimental design module, we construct a comparison experiment against the three baselines introduced in Section 5.1. We employ the search module and Literature Processing module of BioResearcher for 15 research objectives listed in Appendix C (Table C3), thereby obtaining reports and analyses that serve as a knowledge base. The three baselines can invoke search tools to retrieve relevant content from this knowledge base. All experiments are repeated three times, and the evaluation results are averaged and presented in Table 6.
@@ 6 Conclusion
[…]
Yeyun Gong and Chen Lin are the corresponding authors. Chen Lin is supported by the Natural Science Foundation of China (No. 62432011) and Zhongguancun Academy (No. 20240310).
[…]
Appendix A-F.
[…]
If you find our work useful in your research, please cite us using the following BibTeX:
@article{BioResearcher,
author = ”Yi Luo and Linghang Shi and Yihao Li and Aobo Zhuang and Yeyun Gong and Ling Liu and Chen Lin”,
title = ”From Intention To Implementation: Automating Biomedical Research via LLMs”,
journal = ”SCIENCE CHINA Information Sciences”,
year = ”2025”
Appendix A Limitations and Future Works
Limitations
In this paper, BioResearcher effectively automates the entire biomedical research process, from literature and dataset searches to experimental design and execution, given a research objective and specified conditions. However, several limitations warrant discussion. First, the system does not achieve complete success in executing experiments without manual intervention. This limitation is partly due to the need for further enhancement of the code generator agent’s performance. Second, certain anomalies, such as the unavailability of resources specified in the experimental protocols, also hinder fully automated execution. It highlights the necessity of anticipating a broader range of exceptional scenarios and developing corresponding solutions in future work. Third, during our practice, we observed that excessively long inputs often result in LLMs producing shorter, more generalized outputs. To enable LLMs to generate detailed and comprehensive experimental protocols, we employed the multi-step, section-wise processing approach at various stages, including experimental report generation and experimental design. However, this iterative method increases the overall cost. Fourth, the framework currently does not formally model the research process as a constrained optimization problem. This limits its ability to optimally balance competing objectives (e.g., accuracy vs. computational cost) or dynamically allocate resources (e.g., lab equipment, cloud computing budgets), particularly in complex, multi-agent research scenarios.
Future Works
To further enhance and expand the capabilities of BioResearcher, we propose four key future research directions. First, BioResearcher currently supports only dry lab experiments. Future research could explore integrating automated wet lab technologies, such as Cloud Labs [70], to extend the system’s applicability to wet lab studies. Additionally, its utility could be expanded to other scientific disciplines through three modular adaptations: (1) Extensible code modules supporting additional languages (Python/Julia) and domain-specific tools (LabVIEW for physics experiments); (2) Customizable knowledge bases using field-specific repositories (arXiv for CS/physics, ChemRxiv for chemistry); (3) Specialized innovation modules employing generative AI for hypothesis generation in creativity-driven domains. Second, hallucination remains a significant challenge in LLM-based applications, where the model may generate factually incorrect content. This issue is particularly critical in scientific research due to the model’s limited domain-specific knowledge. Addressing LLM hallucination continues to be an essential area for future work. Third, incorporating human oversight mechanisms is crucial for ensuring reliability and ethical compliance in complex scientific workflows. We propose three avenues for integrating human intervention: (1) Confidence-Driven Human Intervention, where a confidence threshold mechanism prompts human validation when the reviewer agent’s confidence falls below a predefined threshold; (2) Iterative Refinement via Natural Language Feedback, allowing users to provide corrections (e.g., missing controls or outdated references), which the AI agents will process and update accordingly while maintaining audit trails; and (3) Ethical Decision-Making Protocol, introducing an automated ethics checkpoint that flags high-risk proposals, mandates human approval for sensitive experiments, and integrates institutional ethics guidelines into the system’s constraints. Future work will quantitatively evaluate the trade-offs between automation and human intervention to optimize system performance and usability. Fourth, incorporating constrained optimization techniques would enable resource-aware scheduling algorithms that balance computational budgets, equipment availability, and temporal constraints through multi-objective optimization – particularly valuable for large-scale collaborative studies.
Appendix B Details of Baselines Implementations
@@
We implement all baselines using the AutoGen framework [71] and adopt OpenAI’s text-embedding-ada-002 as the embedding model in RAG applications.
Appendix C Details of Research Objectives
@@
Appendix D Scoring Criteria
In this section, we outline the specific scoring criteria used in our experiments. Table 11 presents the criteria for evaluating a paper’s helpfulness in achieving a specific research objective. Additionally, Table 12 displays the criteria for assessing the quality of an experimental report across four dimensions.
Table 11: Scoring criteria for the helpfulness of papers.
Rating Description
1 Not Helpful at All:
2 Slightly Helpful:
3 Moderately Helpful:
4 Very Helpful:
5 Extremely Helpful:
Table 12: Scoring criteria for experimental reports.
Dimension Criterion
Appendix E Cost Analysis
@@
Appendix F Comparison of protocols generated by 4 systems.
@@
Table 15 presents the evaluation results of the four assessment models on the protocols described in Section 5.2. As shown in the table, our method consistently outperforms the baseline systems across all four evaluation models.
Table 15: Quality of experimental protocols generated by various systems evaluated by different models.
Evaluator Method Completeness Detail Correctness Logical Soundness Structure Overall $l_{steps}$ $n_{total~{}steps}$
GPT-4o RAG 0.392 0.765 0.470 0.991 0.952 3.570 1.286 9.167
ReAct 0.361 0.617 0.487 0.972 0.950 3.387 1.237 5.792
Plan and Execute 0.366 0.617 0.487 1.000 0.935 3.405 1.194 6.375
[…]
o3-mini-3-31 RAG 0.376 0.731 0.488 0.960 0.967 3.522 1.286 9.167
ReAct 0.329 0.608 0.479 0.958 0.956 3.330 1.237 5.792
Plan and Execute 0.375 0.621 0.479 0.943 0.960 3.378 1.194 6.375
[…]
gemini-2.0-flash RAG 0.397 0.567 0.488 0.982 0.820 3.254 1.286 9.167
ReAct 0.329 0.455 0.490 0.966 0.795 3.036 1.237 5.792
Plan and Execute 0.341 0.475 0.488 0.971 0.827 3.102 1.194 6.375
[…]
deepseek-V3 RAG 0.454 0.685 0.486 0.960 0.892 3.476 1.286 9.167
ReAct 0.435 0.629 0.478 0.957 0.885 3.385 1.237 5.792
Plan and Execute 0.437 0.633 0.476 0.945 0.879 3.370 1.194 6.375
[…]
Average RAG 0.405 0.687 0.483 0.973 0.908 3.456 1.286 9.167
ReAct 0.364 0.577 0.484 0.963 0.897 3.285 1.237 5.792
Plan and Execute 0.380 0.587 0.483 0.965 0.900 3.314 1.194 6.375
BioResearcher $\textbf{0.659}_{{\uparrow\textbf{0.254}}}$ $\textbf{0.893}_{{\uparrow\textbf{0.206}}}$ $\textbf{0.895}_{{\uparrow\textbf{0.411}}}$ $0.953_{\downarrow 0.020}$ $0.891_{\downarrow 0.017}$ $\textbf{4.292}_{{\uparrow\textbf{0.836}}}$ 7.327 33.958